Determine wheather the following sets are subspaces of R

3

{(a,b,c) : a

2+ b

2+ c

2 ≤1, a,b,c ∈R}


1
Expert's answer
2021-12-23T17:57:21-0500

Let us determine wheather the set "W=\{(a,b,c) : a^2+ b^2+ c^2 ≤1, a,b,c ∈\R\}" is a subspace of "\R^3." Taking into account that "1^2+0^2+0^2=1\le1," we conclude that "(1,0,0)\in W." On the other hand, "(1,0,0)+(1,0,0)=(2,0,0)\notin W" because of "2^2+0^2+0^2=4\ge1." Therefore, "W" is not a subspace of "\R^3."